OpenStudy (anonymous):

Simplify ____ √-32 A.)4i√2 B.)-6i√5 C.)-6√5 D.)6√5

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

and please justifie your answer

OpenStudy (anonymous):

ok what happens to 6 if you put it back under the radical?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

ok if we take 6^2 that is going to = the same as \[\sqrt{36}\]

OpenStudy (anonymous):

6

OpenStudy (anonymous):

now if we take 36 and multiply it by the 5 we left under the radical we are going to get 180

OpenStudy (anonymous):

so this already proves b wrong. correct? look at the negitive sign under the radical when it comes out we indicate that with putting a "i" next to our answer and the negitive sign disapears so we now know the answer is?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

explain to me how you came up with that.

OpenStudy (anonymous):

i know we had 6 and we drop the - sigh so it has to be D

OpenStudy (anonymous):

what is 6^2 times 5?

OpenStudy (anonymous):

36*5=180 so we know it's not that one

OpenStudy (anonymous):

180

OpenStudy (anonymous):

so it must be A

OpenStudy (anonymous):

the correct answer is A do you see how it works out?

OpenStudy (anonymous):

besides the trail and error methode we just worked out?

OpenStudy (anonymous):

yes thanks so much

OpenStudy (anonymous):

no problem good luck to you